• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 16:02:45 on April11,2008

    Entry number 231401

    keyword=day shift summary

    Day Shift Summary
    Pi0 production kinematics D

    B. Wojtsekhowski
    G. Jin

    8:00-9:27 kinematics D, run # 2784-2791
    9:28- no beam

    10:30 BB HV alarm 14.6 trip channels from s1 to s7 trip
    11:15 control access
    12:12-13:41 ps3=1, ps7=1, run on cosmics, test for bob, rum # 2795
    13:43- ps3=1, ps7=10000, test for bob, run # 2796-99
    +ROC5 died when beam tuning started
    15:50 restart production with PS5=PS6=1 per BM
    run 2800 - production in progress